OpenStudy (anonymous):

14r + 43 = 25r + 219 R = ?

OpenStudy (anonymous):

@31946 please help

OpenStudy (anonymous):

r=-16

OpenStudy (anonymous):

combine the r terms on one side and the others on the other side 14r -25r = 219 -43 -11r = 176 divide both sides by -11 to get r